In 1993, Waterway Company completed the construction of a building at a cost of $2,380,000 and first occupied it in January 1994. It was estimated that the building will have a useful life of 40 years and a salvage value of $71,200 at the end of that time.

Early in 2004, an addition to the building was constructed at a cost of $595,000. At that time, it was estimated that the remaining life of the building would be, as originally estimated, an additional 30 years, and that the addition would have a life of 30 years and a salvage value of $23,800.

In 2022, it is determined that the probable life of the building and addition will extend to the end of 2053, or 20 years beyond the original estimate.

Using the straight line method annual depreciation from 1994 to 2003 is $57,720 a year.

Annual depreciation from 2004 to 2021 is $76,760 per year.

Compute the annual depreciation to be charged beginning with 2022. Answer is $28,785 I cannot figure out how they arrived at this answer.
Answer #1
Cost 2380000
Less: Salvage value 71200
Depreciable cost 2308800
Useful life 40
Annual depreciation from 1994 through 2003 57720
Cost 2380000
Add: Additions 595000
Total cost 2975000
Less: Accumulated depreciation from 1994 through 2003 577200 =57720*10
Book value, Jan 2004 2397800
Less: Salvage value 95000 =71200+23800
Depreciable cost 2302800
Useful life 30
Annual depreciation from 2004 through 2021 76760
Book value, Jan 2004 2397800
Less: Accumulated depreciation from 2004 through 2021 1381680 =76760*18
Book value, Jan 2022 1016120
Less: Salvage value 95000 =71200+23800
Depreciable cost 921120
Divide by Remaining Useful life 32 =50-18
Annual depreciation expense—building 28785